OpenStudy (anonymous):

I need help with 3 x (1/2)^x an exponential function, can someone go through domain, range, horozontal aysmptote, and graphing

OpenStudy (anonymous):

the range is y>1

OpenStudy (anonymous):

if x is 0 y=2 the domain is all real numbers.

OpenStudy (cwrw238):

Domain is all Real x

OpenStudy (anonymous):

and the horizontal asymptote is 1

OpenStudy (michele_laino):

please keep in mind that the graph of y=f(x)=2^x, is:  so in order to get your graph we have to add to y-corrdinates of all point of 2^x, the quantity +1

OpenStudy (anonymous):

I don't understand the drawing of the graph

OpenStudy (michele_laino):

it is a standard graph which you can find on your textbook

OpenStudy (cwrw238):

do you understand why it pass through y = 1?

OpenStudy (michele_laino):

for example, I write these limits: \[\lim _{x \rightarrow +\infty} 2^{x}=+\infty\] and \[\lim _{x \rightarrow -\infty}2^{x}=0\] do you agree?

OpenStudy (cwrw238):

ooops sorry it should pass through y=2

OpenStudy (anonymous):

I'm sorry but I have the wrong problem up, it's \[3 \times (1/2)^{x}\]

OpenStudy (anonymous):

yep. make x 0 and you get 1 + 1.

OpenStudy (anonymous):

the y intercept is at y=3

OpenStudy (cwrw238):